# Peter Reinhard Hansen§    and   Chen Tong

### v3

@@ 3.2 Moments in Autoregressive Gamma Model
From the estimated model we compute the term structure of conditional moments, $\mathbb{E}_{T}[X_{T+H}^{r}]$. The first conditional moment, $r=1$, is given directly from (10), and we will apply Theorem 2 to obtain other moments. The realized variance measures the second moment of returns, such that $r=\frac{1}{2}$ corresponds to the conditional volatility (standard deviation of returns). Similarly, $r=\frac{3}{2}$ and $r=2$ measures of the conditional skewness (of the absolute value) and the conditional kurtosis, respectively. The inverse of volatility corresponds to the negative moment, $r=-\frac{1}{2}$. This moment is interesting because it is used for Shape ratio forecasting and asset allocation, where the inverse of covariance matrix is employed. We compute these four moments by Theorem 2. The analytical form of the conditional MGF of $X_{T}$ is given in Proposition 2.
@@
[…]
@@ Appendix A Appendix of Proofs
Some notation: We write +=$\mathbb{C}^{+}=\{z\in\mathbb{C}:\mathrm{Re}[z]>0\}$ to denote the complex numbers with positive real part.
In the proof of Theorem 1, we will need the following result (the Laplace transformation of $x^{r}$).
@@
Let $r+1\in\mathbb{C}^{+}$ and $z\in\mathbb{C}^{+}$. Then
$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=\frac{\Gamma(r+1)}{z^{r+1}}.$ (A.1)
@@
Consider the case where $z$ is real and positive, i.e. $z=s>0$. From the definition of the gamma function and the substitution, $y\mapsto sx$, we have $\Gamma(r+1)=\int_{0}^{\infty}y^{r}e^{-y}\mathrm{d}y=\int_{0}^{\infty}s^{r}x^{r}e^{-sx}s\mathrm{d}x$, which proves the identity for any $z\in\mathbb{R}^{+}$. Both sides of (A.1) are holomorphic functions for $z\in\mathbb{C}^{+}$, and since they agree on $z\in\mathbb{R}^{+}$, it follows by the identity theorem that they agree for all $z\in\mathbb{C}^{+}$. [^9] We thank Christian Berg for suggesting this simplification of the proof. A different, direct proof for all $z\in\mathbb{C}^{+}$, is presented in the Supplementary Material. ∎
Lemma A.2.
For $x<0$, $\operatorname{Re}(z)>0$, and $\operatorname{Re}(r)>-1$ we have
$\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t=0.$
For $x=0$, the identity holds for $\operatorname{Re}(r)>0$.
Proof.
We first consider the case $x=0$. The result follows directly from,
$\int_{-\infty}^{+\infty}\frac{e^{z\cdot 0}}{z^{r+1}}\mathrm{d}t=\int_{-\infty}^{+\infty}\frac{1}{(s+it)^{r+1}}\mathrm{d}t=\left.\frac{i}{r}\frac{1}{(s+it)^{r}}\right|_{-\infty}^{+\infty},$
because the last term is zero for ${\rm Re}\left(r\right)>0$.
Next, for $x<0$ we have
$\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t$ $=\int_{-\infty}^{+\infty}\frac{e^{\left(s+it\right)x}}{\left(s+it\right)^{r+1}}\mathrm{d}t$
@@
$=-\int_{+\infty}^{-\infty}\frac{e^{\left(s-i\tau\right)x}}{\left(s-i\tau\right)^{r+1}}\mathrm{d}\tau,\quad\tau=-t$
@@
$=\int_{-\infty}^{+\infty}\frac{e^{\left(s-i\tau\right)x}}{\left(s-i\tau\right)^{r+1}}\mathrm{d}\tau$
$=e^{sx}\int_{-\infty}^{+\infty}\frac{e^{-i\tau x}}{\left(s-i\tau\right)^{r+1}}\mathrm{d}\tau$
$=\lim_{R\rightarrow\infty}e^{sx}\int_{-R}^{+R}f(\tau)\mathrm{d}\tau,$
where $f(\tau)=\frac{e^{-i\tau x}}{\left(s-i\tau\right)^{r+1}}$. The seemingly odd change of variables, $\tau=-t$, serves two important purposes that we explain below.
@@
We will use a closed contour integral in the complex plane with two part. The first part is the line segment along the real axis, from $-R$ to $R$, and the second part is the semicircular arc, $C_{R}=\{Re^{i\theta},\theta\in[0,\pi]\}$. The integral over the line segment, $\int_{-R}^{+R}f(\tau)\mathrm{d}\tau$, is the object of interest, whereas the semicircle defines an auxiliary integral that is used to close the contour, such that we can apply the Cauchy Integral Theorem. The first reason for the change of variable, $\tau=-t$, is to place the pole of $f(\zeta)$ outside the closed contour for all $R>0$ (the pole is at $\zeta=-is$), such that $0=\int_{-R}^{R}f(\zeta)\mathrm{d}\zeta+\int_{C_{R}}f(\zeta)\mathrm{d}\zeta$ by the Cauchy Integral Theorem. For the integral along the arc, we have the expression,
$\int_{C_{R}}f(\zeta)\mathrm{d}\zeta=\int_{C_{R}}e^{-i\zeta x}g(\zeta)\mathrm{d}\zeta=\int_{C_{R}}e^{ia\zeta}g(\zeta)\mathrm{d}\zeta,$
@@
where $a=-x>0$ and $g(\zeta)=\frac{1}{(s-i\zeta)^{r+1}}$ is continuous and satisfies $\lim_{|\zeta|\rightarrow\infty}g(\zeta)=0$ for all $r+1\in\mathbb{C}^{+}$. The second reason for the change of variable is that it enables us to apply Jordan’s lemma, which requires $a>0$ in the expression above. By Jordan’s lemma we have that $\lim_{R\rightarrow\infty}\int_{C_{R}}e^{-ix\zeta}g(\zeta)\mathrm{d}\zeta=0$, which implies that
$\int_{-\infty}^{+\infty}f(\tau)\mathrm{d}\tau=\lim_{R\rightarrow\infty}\int_{-R}^{+R}f(\tau)\mathrm{d}\tau=-\lim_{R\rightarrow\infty}\int_{C_{R}}f(\zeta)\mathrm{d}\zeta=0.$
This completes the proof. ∎
@@
Proof of Lemma 1. Consider first the case $x\geq 0$. Since $\mathrm{Re}[z]=s>0$, we have the following identity from Lemma A.1,
$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=$ $\frac{\Gamma(r+1)}{z^{r+1}},\qquad\text{for all}\quad\mathrm{Re}[r]>-1.$
This is the Laplace transform of $x^{r}$ and the inverse Laplace transform is
$x^{r}=\frac{1}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{\Gamma(r+1)}{z^{r+1}}e^{zx}\mathrm{d}z,\quad x>0,$
provided the integral is well defined. Using $x>0$, we can rewrite the integral as
$|x|^{r}=x^{r}=\frac{\Gamma(r+1)}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}z=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,$ (A.2)
@@
which is well-defined for all $\mathrm{Re}[r]>-1$ (because $x>0$). For $x=0$ the expression is well-defined for all $\mathrm{Re}[r]>0$. This completes the proof of the third expression in Lemma 1. We proceed with
$|x|^{r}$ $=x^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t+\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t$
$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\qquad\quad x\geq 0,$ (A.3)
where the second term is zero by Lemma A.2.
@@
$|x|^{r}=\left(-x\right)^{r}$ $=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad x<0.$
@@
Adding $\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t$ to the right-hand side, which equals zero Lemma A.2, we arrive at
$|x|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}+e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\text{for }x<0.$ (A.4)
Combining (A.3) and (A.4) proves the first expression in Lemma 1.
From (A.2), we also have
$x^{k}$ $=\qquad\quad\frac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{+zx}}{z^{k+1}}\mathrm{d}t,\quad x>0,$
@@
$x^{k}$ $=\left(-1\right)^{k}\frac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{k+1}}\mathrm{d}t,\quad x<0.$
@@
This, and $\Gamma(k+1)=k!$ for $k\in\mathbb{N}$, proves the second expression in Lemma 1. This completes the proof. $\square$
Proof of Theorem 1. The proof is the same as in the proof the Theorem 2 by using the first formula in Lemma 1. $\square$
Proof of Theorem 2. We prove the result for discrete and continuous distributions separately. Consider first a non-negative discrete random variable, $X$, with support $\mathcal{X}$. Rewriting $\mathbb{E}[X^{r}]=\sum_{x\in\mathcal{X}}x^{r}p(x)$ using the third formula in Lemma 1 yields
$\mathbb{E}[X^{r}]$ $=\frac{\Gamma(r+1)}{2\pi}\sum_{x\in\mathcal{X}}\left[\int_{-\infty}^{\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t\right]p(x)$
$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{\infty}\sum_{x\in\mathcal{X}}\frac{e^{zx}}{z^{r+1}}p(x)\mathrm{d}t$
$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{\infty}\frac{M_{X}\left(z\right)}{z^{r+1}}\mathrm{d}t,$
for $r+1\in\mathbb{C}^{+}$ if $\Pr(X=0)=0$ and $r\in\mathbb{C}^{+}$ if $\Pr(X=0)>0$. For a continuous distribution, the same line of arguments gives us,

### v4

@@ 3.2 Moments in Autoregressive Gamma Model
From the estimated model we compute the term structure of conditional moments, $\mathbb{E}_{T}[X_{T+H}^{r}]$. The first conditional moment, $r=1$, is given directly from (15), and we will apply Theorem 1 to obtain other moments. The realized variance measures the second moment of returns, such that $r=\frac{1}{2}$ corresponds to the conditional volatility (standard deviation of returns). Similarly, $r=\frac{3}{2}$ and $r=2$ measures of the conditional skewness (of the absolute value) and the conditional kurtosis, respectively. The inverse of volatility corresponds to the negative moment, $r=-\frac{1}{2}$. This moment is interesting because it is used for Sharpe ratio forecasting and asset allocation, where the inverse of covariance matrix is employed. We compute these four moments by Theorem 1. The analytical form of the conditional MGF of $X_{T}$ is given in Proposition 2.
@@
$\tilde{\beta}_{d}$ $\tilde{\beta}_{w}$ $\tilde{\beta}_{m}$ $\eta$ $\delta$ $\ell$
@@ Appendix A Appendix of Proofs
For $x>0$ and $\operatorname{Re}(r)\in(-1,0)$
$\int_{0}^{\infty}\frac{e^{\pm itx}}{t^{r+1}}\mathrm{d}t=x^{r}\Gamma(-r)e^{\mp\frac{i\pi r}{2}}.$
@@
From Gradshteyn and Ryzhik, (2007, 3.761.4 and 3.761.9) we have
@@
$\int_{0}^{\infty}t^{\mu-1}\cos(tx)\mathrm{d}t$ $=$ $\frac{\Gamma(\mu)}{x^{\mu}}\cos\left(\frac{\pi\mu}{2}\right),\quad x>0,\operatorname{Re}(\mu)\in(0,1),$
@@
$\int_{0}^{\infty}t^{\mu-1}\sin(tx)\mathrm{d}t$ $=$ $\frac{\Gamma(\mu)}{x^{\mu}}\sin\left(\frac{\pi\mu}{2}\right),\quad x>0,\left|\operatorname{Re}(\mu)\right|\in(0,1).$
Using $e^{\pm itx}=\cos(tx)\pm i\sin(tx)$, we have
$\int_{0}^{\infty}t^{\mu-1}e^{\pm itx}\mathrm{d}t$ $=$ $\int_{0}^{\infty}t^{\mu-1}[\cos(tx)\pm i\sin(tx)]\mathrm{d}t$
$=$ $\frac{\Gamma(\mu)}{x^{\mu}}\cos\left(\frac{\pi\mu}{2}\right)\pm i\frac{\Gamma(\mu)}{x^{\mu}}\sin\left(\frac{\pi\mu}{2}\right)=\frac{\Gamma(\mu)}{x^{\mu}}e^{\pm\frac{i\pi\mu}{2}}.$
The result now follows by substituting $\mu=-r$, and noting that $\operatorname{Re}(\mu)\in(0,1)$ for $\operatorname{Re}(r)\in(-1,0)$. ∎
Proof of Lemma 1. Observe that
$\int_{-\infty}^{+\infty}\frac{e^{-itx}}{(s+it)^{\nu}}\mathrm{d}t=\int_{-\infty}^{+\infty}\frac{e^{i\tau x}}{(s-i\tau)^{\nu}}\mathrm{d}\tau=\lim_{R\rightarrow\infty}\int_{-R}^{+R}f(\tau)\mathrm{d}\tau,$
where $f(\tau)=\frac{e^{i\tau x}}{(s-i\tau)^{\nu}}$. Here, we applied the change of variables, $\tau=-t$, for two purposes, which will be explained below.
@@
Let $x>0$ and consider a closed contour integral that consists of the line segment along the real axis, from R-R to $R$, and the semicircular arc, $C_{R}=\{Re^{i\theta},\theta\in[0,\pi]\}$. Our object of interest is the integral over the line segment, $\int_{-R}^{+R}f(\tau)\mathrm{d}\tau$, whereas the semicircle defines an auxiliary integral used to close the contour. The first reason for the change of variable, $\tau=-t$, is to ensure that the pole of $f(\zeta)$ (located at $\zeta=-is$) lies outside the closed contour for all $R>0$ because $s>0$. Thus, by the Cauchy Integral Theorem, $0=\int_{-R}^{R}f(\zeta)\mathrm{d}\zeta+\int_{C_{R}}f(\zeta)\mathrm{d}\zeta$. For the integral along the arc, we have the expression,
$\int_{C_{R}}f(\zeta)\mathrm{d}\zeta=\int_{C_{R}}e^{i\zeta x}g(\zeta)\mathrm{d}\zeta=\int_{C_{R}}e^{ia\zeta}g(\zeta)\mathrm{d}\zeta,$
@@
where $a=x>0$ and $g(\zeta)=\frac{1}{(s-i\zeta)^{\nu}}$ is continuous and satisfies $\lim_{|\zeta|\rightarrow\infty}g(\zeta)=0$ for all $\operatorname{Re}(\nu)>0$. The second reason for the change of variable is that it enables us to apply Jordan’s lemma with $a>0$, which gives us that $\lim_{R\rightarrow\infty}\int_{C_{R}}e^{ix\zeta}g(\zeta)\mathrm{d}\zeta=0$. Thus, by Jordan’s lemma and the Cauchy Integral Theorem, the result now follows.
For $x=0$ the result follows directly by
$\int_{-\infty}^{+\infty}\frac{1}{(s+it)^{\nu}}\mathrm{d}t=\left.\frac{i}{\nu-1}\frac{1}{(s+it)^{\nu-1}}\right|_{-\infty}^{+\infty},$
because the last term is zero for $\operatorname{Re}(\nu)>1$. This completes the proof. $\square$
@@
Proof of Lemma 2. Consider first the case $x\geq 0$. From the Laplace transformation of $x^{r}$,
$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=\frac{\Gamma(r+1)}{z^{r+1}},\qquad\operatorname{Re}(r)>-1,\quad\operatorname{Re}(z)>0,$ (A.1)
see e.g. Gradshteyn and Ryzhik, (2007, 3.381.4), and the inverse Laplace transform is $x^{r}=\frac{1}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{\Gamma(r+1)}{z^{r+1}}e^{zx}\mathrm{d}z$, provided the integral is well defined. For $x>0$, we can rewrite the integral as
$|x|^{r}=x^{r}=\tfrac{\Gamma(r+1)}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}z=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,$ (A.2)
@@
which is well-defined for all $\operatorname{Re}(r)>-1$. For $x=0$ the expression is well-defined for $\operatorname{Re}(r)>0$. This completes the proof of the second expression in Lemma 2. Because $\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t=0$ for $x>0$ by Lemma 1, we have
$|x|^{r}=x^{r}$ $=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t+\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t$
$=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\qquad\quad x>0.$ (A.3)
@@
$|x|^{r}=\left(-x\right)^{r}$ $=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad x<0.$
@@
Adding $\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t$ to the right-hand side, which equals zero Lemma 1, we arrive at
$|x|^{r}=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}+e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\text{for }x<0.$ (A.4)
Combining (A.3) and (A.4) proves the first expression in Lemma 2 for $x\neq 0$ and $\mathrm{\operatorname{Re}}(r)>-1$. For $x=0$ and $\mathrm{\operatorname{Re}}(r)>0$, the left-hand side is $0^{r}=0$ and the integral is zero by Lemma 1. For $x=r=0$, the left-hand side is $0^{0}=1$ and the integral simplifies as follows,
$\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}+e^{zx}}{z^{r+1}}\mathrm{d}t=\tfrac{1}{2\pi}\int_{-\infty}^{+\infty}\frac{2}{z}\mathrm{d}t=1,$
where we used (2). This completes the proof of first formula.
Finally, the last expression in Lemma 2 follows from
$x^{k}$ $=\qquad\quad\tfrac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{+zx}}{z^{k+1}}\mathrm{d}t,\quad x>0,$
@@
$x^{k}$ $=\left(-1\right)^{k}\tfrac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{k+1}}\mathrm{d}t,\quad x<0,$
@@
and $\Gamma(k+1)=k!$, where we used (A.2) with $k\in\mathbb{N}$ and $x>0$ and x>0-x>0, respectively. This completes the proof. $\square$
Proof of Theorem 1. Set $x=X-\xi$ in the first expression of Lemma 2 and take expected value to both sides, then we get